LOG-IN
Displaying report 1-1 of 1.
IOO (Laser Bench)
Print this report.
stefan.ballmer - 18:26, Friday 08 June 2018 (5227)Get code to link to this report
PMC work today
Keiko, Stefan

- We installed the PMC dampers (see picture). It turns out they are really effective in damping the 5kHz resonance, which I presume is the first full body mode.
- We attached the measured transfer function. Note that we can now easily get 800Hz UGF, with a 1/f integrator down to 3.4Hz.
- We prpperly the PMC trans camera after reflecting of a steering mirror and a beam splitter.
- The transmitter beam from that beam splitter is used for the PMC_TRANS photodiode, which we hooked up to the ADC. Note that this PD is a Thorlabs unit that is still powered by its own AC power supply.
- We also temporarily hooked up the DC output from the LSC REFL diode to the ADC as PMC_REFL. Eventually we will set up an independent Thorlabs PD for that too.
- We started to align the FSS path: We took out the EOM and the first lens for the initial coarse alingment. We stopped there because of other priorities. Still TBD.
- Verified that the SLOW DAC signal into the TTFSS Interface box looks healthy. SO the problem is with the analog electroic. More debugging needed.

Pictures / Files:
1) PMC
2) OLG of PMC locking loop - the 5kHz resonance is almost gone.
3&4) CSV and MAT file from the Mokulab open loop measurement
Images attached to this report
5227_20180608112300_img0442.jpg 5227_20180608112336_mokubodeanalyzerdata20180608172628screenshot1.png
Non-image files attached to this report